OpenStudy (anonymous):

15. Which is larger for a disk of mass M and radius R that is rolling without slipping: its transnational or its rotational kinetic energy? PLZZZ TELL

OpenStudy (vincent-lyon.fr):

KE = 1/2 Mv² + 1/2 I w² with I = MR²/2 and v = wR you can conclude.